OpenStudy (anonymous):

Tracy needs to find the length and width of the equipment storage room where she works. The room is 2 meters longer than it is wide. Its perimeter is 24 meters. Fine the width in meters.

OpenStudy (anonymous):

Call length, l and width,w. We have l=w+2. since the length is 2 meters longer the width. The perimeter is: 2l+2w=24, substituting l=w+2: 2(w+2)+2w=24 solve for w. Once you have w, go back to l=w+2 to get the length, l.

OpenStudy (anonymous):

You lost me... what do I do to figure out the answer?

OpenStudy (anonymous):

solve: 2(w+2)+2w=24

OpenStudy (anonymous):

but i dont know what w is

OpenStudy (anonymous):

you need to be able to solve basic algebraic equations to get the answer. You have to find w, using algebra...

OpenStudy (anonymous):

I'll give you the first step: 2w+4+2w=24 go from there...

OpenStudy (anonymous):

idk how to do that

OpenStudy (anonymous):

I'd review your basic algebra... basically, collect "like" terms (example 3x+10x can be simplified to 13x). isolate unknown variable (example 5x+20=80, becomes 5x=60...then we get x=12). good luck :)

OpenStudy (anonymous):

its 5

OpenStudy (anonymous):

yep. so now we have w=5. and we know that l=w+2. Input w=5 to get the length, l

OpenStudy (anonymous):

actually you don't even need to since the question only asked for the width. so you're done

OpenStudy (anonymous):

ik.

OpenStudy (anonymous):

sg. gichywt.